OpenStudy (anonymous):

PLEASE HELP!!! (sqrt 3 - i) in exponetial form

OpenStudy (anonymous):

\[\sqrt{3}-i\]

OpenStudy (anonymous):

you need two numbers \(r\) and \(\theta\) to write \[\sqrt3-i=r(\cos(\theta)+i\sin(\theta))\] namely \(r\) and \(\theta\)

OpenStudy (anonymous):

i'm so lost. can you help more?

OpenStudy (anonymous):

sure \(r\) is the easiest to find it is \(|a+bi|=\sqrt{a^2+b^2}\) which in our case is \[\sqrt{\sqrt3^2+1^2}=\sqrt{3+1}=\sqrt4=2\]

OpenStudy (anonymous):

then you need \(\theta\)

OpenStudy (anonymous):

\[\cos(\theta)=\frac{a}{4}=\frac{\sqrt3}{2}\] \[\sin(\theta)=\frac{b}{r}=-\frac{1}{2}\] from this you can find \(\theta\)

OpenStudy (anonymous):

theta is the arc tan \[\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so how do i put this into exponential form from there?

OpenStudy (anonymous):

careful here!!

OpenStudy (anonymous):

use sine and cosine

OpenStudy (anonymous):

find the point on the unit circle with the coordinate \[(\frac{\sqrt3}{2},-\frac{1}{2})\]

OpenStudy (anonymous):

yes \[\frac{ 11\pi }{ 6}\]

OpenStudy (anonymous):

but whre does the exponent come in?

OpenStudy (anonymous):

are you trying to write this as \[\large \sqrt3-i=re^{i\theta}\]?

OpenStudy (anonymous):

or as \[\sqrt3-i=r(\cos(\theta)+i\sin(\theta))\]?

OpenStudy (anonymous):

the first one using euler's formula

OpenStudy (anonymous):

ok well it is the same exactly

OpenStudy (anonymous):

the \(r\) is still \(r\) and same \(\theta\)

OpenStudy (anonymous):

you have a choice of angles

OpenStudy (anonymous):

but if you pick \(\frac{11\pi}{6}\) then it is \[\large 2e^{\frac{11\pi}{6}i}\]

OpenStudy (anonymous):

ooooohhhhh i didn't know it was this simple! thank you

OpenStudy (anonymous):

lol simple when you know how right? like most things...

OpenStudy (anonymous):

yw